# Context Engineering for AI Agents in Open-Source Software

### v1

@@ 4.3. Evolution (RQ3)
To answer RQ3, we first screened the commit histories of all 155 AGENTS.md files and found that 77 (50%) of them had not been changed after the initial commit, 36 (23%) only once, and 32 (21%) between two and seven times. For this study, we decided to focus on 10 files (6%) with at least 10 commits, which yielded a sample of 169 commits to annotate (37% of the 453 collected commits). We were primarily interested in understanding the types of changes that developers make to AI configuration files. The considered time spans vary between repositories, with some histories spanning a short period with many changes (e.g., neomjs/neo with 49 changes over 19 days) and others spanning longer periods with fewer changes (e.g., gofiber/fiber: 11 changes, 148 days). Although we did not analyze files with fewer than 10 commits in detail, we noticed that the history of these files varies significantly as well, ranging from 0 to 127 days, with 2 to 8 commits. A more detailed analysis of evolution patterns is an important direction for future work.
To understand what developers change in AI configuration files and to inductively identify general categories of changes, we manually reviewed the commits, including the source code diff, the commit messages, and any related issues or pull requests. Two authors developed an initial coding guide and then iteratively refined the emerging categories and descriptions For the analysis, we treated each commit as an isolated change. However, we observed that one change in the commit history can exemplify multiple types of changes rather than only one, but most commits (111, 66%) represented only one change category. Table 2 shows that some change categories such as ‘Add section(s)’ refer to the overall structure of AGENTS.md files and indicate that the change to instructions was rather significant, while others such as ‘Add instruction(s)’ refer to the content of sections and the extent and scope of change was more focused and local. Note that, at this stage, we do not differentiate between single and multiple changes of a category within a file. For example, ‘Add section(s)’ can mean that one section was added or multiple sections were added. Furthermore, we do not discuss the intent of a change. For example, for the category ‘Remove section(s)’ we do not analyze why this may have happened. An analysis of intent is subject of future work.
Table 1. Categories of information provided in AGENTS.md files (last column: number of level 1 or 2 headings).
@@
Conventions/best practices Outlines coding standards, naming/formatting conventions, and best practices for writing consistent and maintainable code. 50
@@
Architecture and project structure Describes how the project or repository is organized, including key directories, modules, components, and relationships between them. 47
@@
Project description Summarizes what the project or agent does, its goals or purposes, and high-level functionality or capabilities. 32
Testing instructions Explains how to execute test suites or individual tests, including tools, commands, and environments. 32
Document metadata Contains file metadata or configuration (e.g., tags). 29
Testing strategy and guidelines Describes the overall approach to testing (unit, integration, end-to-end), test organization, or principles guiding test coverage and design. 24
Tech stack and dependencies Lists programming languages, libraries, frameworks, or other dependencies used in the project. 15
Getting started and setup Covers installation prerequisites, environment setup, and initial steps required to run/use the project locally. 11
References and cheat-sheets Provides a concise list of frequently used commands, API references, or quick tips for developers or users. 9
@@
Patterns and examples Shows reusable patterns, sample agent configs, or example use cases to guide understanding or extensions. 8
@@
Add instruction(s) Add one or more lines of instructions to existing sections. 78
Modify instruction(s) Modify one or more lines of instructions within a section (not related to fixing typos or adding external references). 59
Add section(s) Add new sections to the AGENTS.md file. 26
Remove instruction(s) Remove one or more lines of instructions from existing sections. 23
Modify headings Modify existing section heading title or level. 23
Modify text (minor) Minor changes to content of AGENTS.md file, such as fixing typos. 19
Reformat instruction style Changing visual appearance of content in AGENTS.md file (not related to structure). 10
Remove section(s) Remove sections from AGENTS.md. file 2
@@
Table 2 shows that the most frequent change categories are ‘Add instruction(s)’ and ‘Modify instruction(s)’. For all examined AGENTS.md files, these categories occurred as the first or second change in the history of changes. Looking at commit messages related to changes, we found a few interesting cases. For example, for AGENTS.md in rsyslog, one of the commit messages states AI support: Agent shall no longer call stylecheck.sh. The related change category is ‘Remove instruction(s)’, which indicates that an agent should not perform a certain operation. A commit in eclipse-rdf4j/rdf4j fixed a flaky test and at the same time updated the AGENTS.md file to handle potentially flaky tests during test execution. Another bug fix for the same file fixed a null pointer exception, again adjusting to test execution instructions to prevent related issues. This co-changes of source code, tests, and AI configuration files are a promising direction for future work. In summary, the evolution of the AGENTS.md files in our sample varies, and we did not identify clear patterns in terms of when and how often changes occur. However, we did identify frequent change categories. Based on these categories, it appears that changes early on in the lifetime of AI configuration files are mostly about fine-tuning and adjusting instructions rather than major changes.
@@ 5. Conclusion
In addition to README files written for human contributors, projects increasingly include AI configuration files that provide instructions and context for AI coding agents. In other words, software developers are now writing and maintaining documentation for machines. Our results show that conventions for this new form of documentation are still in flux. Projects differ widely in what they encode (e.g., architecture, conventions, workflows) and how they express it (e.g., prescriptive, prohibitive, explanatory). These stylistic variations mirror broader prompt writing practices and raise questions about whether and how tone and phrasing influence agent adherence or output quality. In this sense, open-source repositories serve as natural laboratories for studying how developers experiment with “talking” to agent-based AI tools.
AI configuration files should be treated as maintained software artifacts: versioned, reviewed, quality-assured, and tested. Future work needs to evaluate how their content, structure, and style affect agent behavior and task performance, and how automated feedback loops could update or refine these files based on observed results. Research could also investigate the co-evolution of source code and related AI configuration files, similar to the co-evolution of source code and comments (Fluri et al., 2009).
Design questions include whether standard schemas could improve interoperability, whether repositories should maintain one centralized file or multiple module-level ones, and how to coordinate instructions for multiple agents. Beyond technical considerations, this new form of documentation has the potential to reshape communication, review, and collaboration patterns in software teams as instructions move from being written for humans to being negotiated between humans and AI. The systematic study of AI configuration files has great potential to provide actionable recommendations to practitioners.

### v4

@@ 4.3. Evolution (RQ3)
To answer RQ3, we analyzed the commit histories of all 155 AGENTS.md files and found that 77 (50%) of them had not been changed, 36 (23%) only once, and 32 (21%) between two and seven times. For this study, we were primarily interested in understanding the types of changes developers make in AI context files. We decided to focus on the 10 files (6%) with at least 10 commits, which yielded a sample of 169 commits to annotate (37% of all collected commits). The resulting modification patterns varied per file, with some histories spanning a short period with many changes (e.g., neomjs/neo: 49 changes over 19 days) and others spanning longer periods with fewer changes (e.g., gofiber/fiber: 11 changes, 148 days). Although we did not analyze files with fewer than 10 commits in detail, we noticed that the history of these files varied significantly as well, ranging from 0 to 127 days, with 2 to 8 commits. A more detailed analysis of evolution patterns is an important direction for future work.
To understand what developers change in AI context files and to inductively identify general change categories, we manually reviewed the commits, including the source code diff, the commit messages, and any related issues or pull requests. Two authors developed an initial coding guide and then iteratively refined the emerging categories and descriptions. We considered each commit as an isolated change. Although most commits (111 commits, i.e., 66%) represented only one change category, we observed commits that represented multiple. Table 2 shows that some categories refer to the overall structure of AGENTS.md files while others refer to the content of specific sections, indicating a narrower scope. Note that we currently do not quantify changes per category, meaning that, e.g., ‘Add section(s)’ can represent a change that added one or multiple sections. Furthermore, we did not label the intent of a change, e.g., why a certain section was added or removed. An analysis of the intent of changes is the subject of our future work.
Table 1. Categories of information provided in AGENTS.md files (last column: number of level 1 and 2 headings).
@@
Conventions Outlines coding standards, naming/formatting conventions, and best practices for writing consistent and maintainable code. 50
@@
Architecture/ structure Describes how the project or repository is organized, including key directories, modules, components, and relationships between them. 47
@@
Goals/purposes Summarizes what the project or agent does, its goals or purposes, and high-level functionality or capabilities. 32
Test execution Explains how to execute test suites or individual tests, including tools, commands, and environments. 32
Metadata Contains file metadata or configuration (e.g., tags). 29
Test strategy Describes the overall approach to testing (unit, integration, end-to-end), test organization, or principles guiding test coverage and design. 24
Tech stack Lists programming languages, libraries, frameworks, or other dependencies used in the project. 15
Setup Covers installation prerequisites, environment setup, and initial steps required to run/use the project locally. 11
References Provides a concise list of frequently used commands, API references, or quick tips for developers or users. 9
@@
Patterns/examples Shows reusable patterns, sample agent configs, or example use cases to guide understanding or extensions. 8
@@
Add instruction(s) Add instruction line(s) to existing sections. 78
Modify instruction(s) Modify instruction line(s) within a section (ignoring typo fixes and references additions). 59
Add section(s) Add new section(s) to the AGENTS.md file. 26
Remove instruction(s) Remove line(s) with instructions from existing sections. 23
Modify heading(s) Modify existing section heading title or level. 23
Modify text Minor changes to content of AGENTS.md file, such as fixing typos. 19
Reformat style Changing visual appearance of content in AGENTS.md file (not related to structure). 10
Remove section(s) Remove sections from AGENTS.md file. 2
@@
Table 2 shows that the most frequent change categories are ‘Add instruction(s)’ and ‘Modify instruction(s)’. For all examined AGENTS.md files, these categories occurred as the first or second change in the history of changes. Looking at commit messages related to changes, we found a few interesting cases. For example, for AGENTS.md in rsyslog, one of the commit messages states AI support: Agent shall no longer call stylecheck.sh. The related change category is ‘Remove instruction(s)’, because the change deleted an instruction. A commit in eclipse-rdf4j/rdf4j fixed a flaky test and updated the AGENTS.md file to handle flaky tests during test execution. Analyzing co-changes of AI context files and other artifacts is a promising direction for future work.
In summary, the evolution of the AGENTS.md files in our sample varies, and we did not identify clear patterns in terms of when and how often changes occur. However, we did identify common change categories. Based on these categories, it appears that changes are mostly made to fine-tune and adjust instructions.
@@ 5. Conclusion
In addition to README files for humans, OSS projects increasingly include AI context files for AI coding agents. In other words, software developers are now writing and maintaining documentation for machines. Our results show that conventions for this new software artifact are still in flux. Projects differ widely in what they encode (e.g., conventions, architecture) and how they express it (e.g., prescriptive vs. prohibitive). These stylistic variations mirror different prompt writing practices. Thus, OSS repositories serve as natural laboratories for studying how developers experiment with “talking” to agent-based AI tools.
AI context files are maintained software artifacts. They are versioned, reviewed, quality-assured, and tested. Future work needs to evaluate how their content, structure, and style affect agent behavior and task performance, and how automated feedback loops could update or refine these files based on observed results. Research should also investigate the co-evolution of source code and related AI context files, similar to the co-evolution of source code and comments (Fluri et al., 2009). Open questions include whether standard schemas could improve interoperability, whether repositories should maintain one or multiple AI context files, and how to coordinate instructions for multiple agents. Beyond technical considerations, this new form of documentation has the potential to reshape communication, review, and collaboration patterns in software teams as instructions move from being written for humans to being negotiated between humans and AI. The systematic study of AI context files has great potential to provide actionable recommendations to practitioners.